Circular No. 7204

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUNGRAZING COMETS D. A. Biesecker, SM&A Corporation and Goddard Space Flight Cen-

ter, reports observations of three more apparent Kreutz sungrazing comets found in images taken with both the C2 and C3 coronagraphs aboard SOHO. Astrometric measurements made by Biesecker (and reduced by B. G. Marsden), together with orbital elements by Marsden, are given on MPECs 1999-M08 (C/1999 K9), 1999-M09 (C/1999 K10), and 1999-M10 (C/1999 L4).

1999 UT  α_{2000}  δ_{2000}  Comet
May 23.698  4^{h}03.^{m}8  +18^{o}13′  C/1999 K9
31.698  4 33.4  +19 47  C/1999 K10
June 1.448  4 33.9  +19 37  C/1999 L4

Comets C/1999 K9 and C/1999 L4 were discovered by A. Vourlidas; C/1999 K10 was discovered by Biesecker. No tail is evident for C/1999 K9 on images obtained during May 23–24. No tail is visible in the May 31–June 1 images of C/1999 K10; this object was extremely faint in the C3 images. A faint tail is visible is visible in the June 1 images of C/1999 L4, which was again very difficult to measure in the C3 frames.

COMET C/1998 M5 (LINEAR)

Ephemeris extension to IAUC 7137:

1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 21 9^{h}20.^{m}61 +24^{o}25.^{′}7 3.129 2.537 46.^{o}3 16.^{o}8 12.5
26 9 25.19 +22 47.6 3.228 2.580 43.0 15.6 12.7
July 1 9 29.74 +21 13.9 3.325 2.623 39.6 14.3 12.8
6 9 34.25 +19 44.2 3.418 2.667 36.3 13.0 12.9
11 9 38.71 +18 17.9 3.508 2.710 32.9 11.7 13.1

Total visual magnitude estimates: May 16.86 UT, 10.9 (M. Reszelski, Szamotuly, Poland, 0.25-m reflector); 24.88, 11.7 (W. Hasubick, Buchloe, Germany, 0.44-m reflector); June 3.89, 12.2 (M. Lehký, Hradec Králové, Czech Republic, 0.42-m reflector); 13.30, 12.5 (N. Biver, Oahu, HI, 0.26-m reflector).

1999 June 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green